A particle has charge -3 nC. A) Find the magnitude and direction of the electric field due to this particle at a point 0.25 m directly above it. B) At what distance from this particle does its electric field have a magnitude of 12 N/C?


1
Expert's answer
2021-05-19T20:41:15-0400

"a) E=kq/r^2, k=9\times 10^9"

"E=-\frac{(9\times 10^9 \times3\times 10^-9)}{0.25^2}=-" "432N\/C"


b) "12=\frac{9\times 10^9 \times 3\times 10^-9}{r^2}"

"r^2=27\/12=2.25"

"r=1.5m"